1. The probability of a teenager owning a surfboard is .43, owning a skateboard is .38, and owning both is .28 If a teenager is selected at random, find the probability that he or she owns a surfboard or a skateboard.

2. In a class consisting of 15 men and 12 women, two homework papers were selected at random. Find the probability that both papers belong to women.

3. An urn contains 6 red balls, 2 green balls, 1 blue ball, and 1 white ball. If a ball is drawn, find the probability of getting a red or white ball.

4. 88% of American children are covered by health insurance. If four children are selected at random, find the probability that none are covered.

5. If a die is rolled 5 times, find the probability of getting at least one 6.

6. Which score has the best relative position?

A) 3.2 where sample mean is 4.6 and s is 1.5

B) 630 where sample mean is 800 and s is 200

C) 43 where sample mean is 50 and s is 5
Answer #1

1)  If a teenager is selected at random, find the probability that he or she owns a surfboard or a skateboard= 0.43+0.38-0.28 = 0.53.

2) the probability that both papers belong to women = 12*11/(27*26) = 0.188

3)  If a ball is drawn, find the probability of getting a red or white ball = 7/10 = 0.7

4) If four children are selected at random, find the probability that none are covered = 0.124 =0.0002